OpenStudy (anonymous):

try and prove that secx tanx(1-sin^2x)=sinx

OpenStudy (badhi):

$$ \sec x \tan x(1-\sin^2x)=\frac{1}{\cos x} \frac{\sin x}{\cos x}(\cos^2 x)$$